OpenStudy (anonymous):

solve the integral: (y-1)/e^y dy

OpenStudy (anonymous):

Ok so I'm about to use integration by parts. int(udv) = uv - int(vdu) u = y-1 du = 1 dy dv = e^-y v = -e^-y This gives: uv - int(vdu) = (y-1)(-e^-y) - int(-e^-y)dy = (y-1)(-e^-y) + int(e^-y)dy (I pulled out a negative from the integral) = (y-1)(-e^-y) - e^-y + C = -y*e^-y + e^-y - e^-y + C = -y*e^-y + C

OpenStudy (anonymous):

THANK YOU!.....i have more if you want to enjoy them!

OpenStudy (anonymous):

You're welcome. Make sure to check the work. Eh... I had to do ~250 integrals last month, so I'm not sure that I would want to do all of them. It's really important that you do them yourself, but if you find a few that you absolutely can't get after several tries, feel free to ask for help.

OpenStudy (anonymous):

i will be asking, thank u very much